OpenStudy (anonymous):

Solve this equation for A: A ÷ 2 = 4

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (cherrilyn):

A = 8

OpenStudy (angela210793):

A/2=4 A=4*2=8